# Catastrophic Goodhart: regularizing RLHF with KL divergence does not mitigate heavy-tailed reward misspecification

### v1

@@ 6.2 Independence assumptions
Theorems 2 and 3 do not require any independence assumption, but Theorems 4, 5, and 6 require that error $X$ and utility $V$ are independent, which seems to be violated in practice. Future work could weaken this assumption, although intuitively obvious ways to weaken it result in the statement being false.  [^4] Suppose that error $X$ is light-tailed conditional on any value of $V$, but our proxy is merely unbiased ($\mathbb{E}[X|V=v]=0$ for all $v$). Then the limit of $V$ under optimization for $X+V$ still depends on the relationship between $X$ and $V$. If they are independent, Theorem 6 says that $\lim_{t\to\infty}\mathbb{E}[V|X+V\geq t]=\infty$. But if $V\sim N(0,1)$, and $X|V\sim N(0,4)$ when $V\in[-1,1]$, otherwise $X=0$, then $\lim_{t\to\infty}\mathbb{E}[V|X+V\geq t]=0$.
@@ 6.4 Relation to previous overoptimization work
Gao et al. (2023) found that optimizing the reward of small reward models causes overoptimization: a decrease in utility with increasing optimization. However, we observed that reward models are light-tailed, and (Theorem 4) that independence combined with light-tailed error prevents overoptimization. We think this discrepancy is explained by dependence between error and utility. Policies optimized for high error may activate features in the proxy reward models that are undesirable according to the true utility function.  [^5] There are other explanations possible. Perhaps better optimization methods would find heavy-tailed reward in open reward models; or OpenAI’s reward models have heavy-tailed error (and their results are straightforwardly explained by our Theorem 1), while open reward models have light-tailed error. More research is needed to understand why high-error completions have low utility and to design reward models that do not suffer from this problem; Perhaps it is possible to construct reward models whose errors are in directions orthogonal to human preferences, so that the large-reward completions do not have lower utility.
@@ A.1 Theorem 1
Intuitively, in a heavy-tailed distribution, events with extremely high $x$ are not very rare, so you don’t pay much of a KL penalty to upweight them so they happen about $1/x$ of the time. This is visually illustrated in Figure 3.
Figure 3: As $t\to\infty$, the mean of $X$ (blue bar) grows without bound while KL divergence $D_{KL}(P_{t}\,\|\,Q)$ (orange bar) goes to 0. The base distribution Q is a Student t-distribution with $df=3$. In this case, high values of X are upweighted to $1/t^{0.8}$; upweighting them to $1/t$ would cause $\mathbb{E}[X]$ to converge to  $1$ while KL divergence goes to zero faster.
@@ A.2 Theorem 2
A.3 Theorem 3
Restatement of Theorem 3.
$\log\frac{P(V)}{Q(V)}=\frac{V-\lambda_{2}}{\lambda_{1}}$ (5)
$P(V)=Q(V)\exp\left(\frac{V-\lambda_{2}}{\lambda_{1}}\right)=Q(V)$ (6)
[…]
@@ A.4 Theorem 4
If $U=X+V$ with $X$ and $V$ both light-tailed, and the distribution of U is continuous, and $\pi^{*}(\beta)\triangleq\arg\max_{\pi}\mathbb{E}[U(\pi)]-\beta D_{KL}(\pi,\pi_{0})$, then $\lim_{\beta\to 0^{+}}\mathbb{E}[V(\pi^{*}(\beta))]=\infty$.
@@
Appendix B Conditioning as alternate model of optimization
Although we think a KL divergence penalty or cap is the most realistic setting for RLHF, it is not the only model of optimization where heavy-tailedness of the error determines whether catastrophic Goodhart occurs. Consider another model of optimization where $U=X+V$ as before, but we simply condition on $U$ being higher than some threshold $t$ [^6] This could model a satisficing agent that takes random acceptable actions. In this case, we are interested in the quantity $\lim_{t\to\infty}\mathbb{E}[V|X+V\geq t]$. If we slightly strengthen the heavy-tailedness and light-tailedness assumptions, heavy-tailed error results in catastrophic Goodhart, while light-tailed error results in arbitrarily high expected utility.
B.1 Conditioning with heavy-tailed error produces catastrophic Goodhart
Theorem 5.
@@ B.1 Conditioning with heavy-tailed error produces catastrophic Goodhart
  • The tail of $V$ is sufficiently lighter than the tail of $X$ such that $\lim_{t\to\infty}\frac{t^{p}\bar{F}_{V}(t)}{\bar{F}_{X}(t)}=0$ for some $p>1$.
@@
In addition to the works cited in the main paper, we make reference to the textbook (Foss et al., 2013) throughout the proof. Many similar results about random variables are present in the textbook.
B.1.1 Proof sketch and intuitions
@@
In Table 1, we abbreviate the condition $X+V>t$ as $c$.
@@
Table 1: A summary of the proof strategy for Theorem 5.
Figure 4: A diagram showing the region boundaries at $-h(t)$, $h(t)$, and $t-h(t)$ in an example where $t=25$ and $h(t)=4$, along with a negative log plot of the relevant distribution:
@@ B.1.3 Lemmas
Lemma 2.19 from (Foss et al., 2013) implies that if $X$ is long-tailed (which it is, because subexponential implies long-tailed), then there is $h(t)$ such that condition (a) holds and $\bar{F}_{X}$ is $h$-insensitive; by Proposition 2.20 we can take $h$ such that $h(t)\leq t/2$ for sufficiently large $t$, implying condition (b). Conditions (c) and (d) follow from being $h$-insensitive. ∎
@@
This is a slight variation on lemma 3.8 from (Foss et al., 2013), and follows from the proof of Lemma 2.37. Lemma 2.37 states that
@@ B.1.5 Bounds on the denominator
B.2 Conditioning with light-tailed error produces arbitrarily high utility
Theorem 8.
@@ B.2 Conditioning with light-tailed error produces arbitrarily high utility
Theorem 8 generalizes a consequence of the "Regressional Goodhart Identity" in (Gao et al., 2023).
@@
Appendix C Additional figures
See figures 5, 6.
@@ Appendix C Additional figures
Figure 5: Histogram and normal probability plot of reward assigned by Pythia RM to random length-1024 token sequences. The Q-Q plot suggests the distribution is approximately normal, which is much lighter-tailed than exponential.
Figure 6: Reward and log-probability for ACG-optimized inputs to Starling 7B-alpha.
Appendix D Hyperparameters for ACG
See table 2.
Table 2: Hyperparameters for ACG
Parameter Value
Context length 133
Iterations 1000
Candidates per seq. position (k) 3
Annealing starting value 9
Annealing ending value 2
Appendix E Assets
@@
We use three models for our experiments: Starling 7B-alpha, Llama 2 7B-chat, and Pythia-1.4B. Starling was developed by Berkeley, and Pythia by EleutherAI. Starling and Pythia models are licensed under Apache-2.0.999https://twitter.com/NexusflowX/status/1770532630645420474  [^10] https://huggingface.co/EleutherAI/pythia-1.4b Llama 2 models were developed by Meta and licensed under a license published by Meta. [^11] https://ai.meta.com/llama/license/

### v2

@@ 5.2 Independence assumptions
Theorems 1-3 do not require any independence assumption, but Theorems 4, 5, and 6 require that error $X$ and utility $V$ are independent, which seems to be violated in practice. Future work could weaken this assumption, although intuitively obvious ways to weaken it result in the statement being false.  [^5] Suppose that error $X$ is light-tailed conditional on any value of $V$, but our proxy is merely unbiased ($\mathbb{E}[X|V=v]=0$ for all $v$). Then the limit of $V$ under optimization for $X+V$ still depends on the relationship between $X$ and $V$. If they are independent, Theorem 6 says that $\lim_{t\to\infty}\mathbb{E}[V|X+V\geq t]=\infty$. But if $V\sim N(0,1)$, and $X|V\sim N(0,4)$ when $V\in[-1,1]$, otherwise $X=0$, then $\lim_{t\to\infty}\mathbb{E}[V|X+V\geq t]=0$.
@@ 5.3 Stronger optimization methods
5.4 Reparameterizing reward
In some cases, a heavy-tailed reward can be reparameterized to make it light-tailed and avoid catastrophic Goodhart; however, in settings where the true reward is heavy-tailed, making reward artificially light-tailed or bounded can result in unintended behavior.
For example, a stock-trading agent should be rewarded by profit, but financial returns are known to be heavy-tailed. If we clip or otherwise transform rewards into a bounded interval, it will have no incentive to take into account huge gains or losses. Since RLHF rewards as implemented in Ziegler et al are unbounded, clipping or transforming rewards could itself cause reward misspecification.
In some cases, e.g. when the reward is not the true intended one, it is possible to reparameterize the reward without adverse effects. In the RL literature for Atari games, rewards are changes in score clipped to $[-1,1]$ (Machado et al., , 2018).
@@ 5.5 Relation to previous overoptimization work
Gao et al., (2023) found that optimizing the reward of small reward models causes overoptimization: a decrease in utility with increasing optimization. However, we observed that reward models are light-tailed, and (Theorem 4) that independence combined with light-tailed error prevents overoptimization. We think this discrepancy is explained by dependence between error and utility. Policies optimized for high error may activate features in the proxy reward models that are undesirable according to the true utility function. [^6] There are other explanations possible. Perhaps better optimization methods would find heavy-tailed reward in open reward models; or OpenAI’s reward models have heavy-tailed error (and their results are straightforwardly explained by our Theorem 1), while open reward models have light-tailed error. More research is needed to understand why high-error completions have low utility and to design reward models that do not suffer from this problem; perhaps it is possible to construct reward models whose errors are in directions orthogonal to human preferences, so that the large-reward completions do not have lower utility.
@@ 6 Conclusion
Acknowledgments and Disclosure of Funding
This work was supported by the Long-Term Future Fund (LTFF). We also thank the anonymous reviewers for their valuable feedback and constructive suggestions.
@@ A.1 Theorem 1
Intuitively, in a heavy-tailed distribution, events with extremely high $x$ are not very rare, so you don’t pay much of a KL penalty to upweight them so they happen about $1/x$ of the time. This is visually illustrated in Figure A.1.
Figure A.1: As $t\to\infty$, the mean of $X$ (blue bar) grows without bound while KL divergence $D_{KL}(P_{t}\,\|\,Q)$ (orange bar) goes to 0. The base distribution Q is a Student t-distribution with $df=3$. In this case, high values of X are upweighted to $1/t^{0.8}$; upweighting them to $1/t$ would cause $\mathbb{E}[X]$ to converge to  $1$ while KL divergence goes to zero faster.
@@ A.2 Theorem 2
$\log\frac{P(V)}{Q(V)}=\frac{V-\lambda_{2}}{\lambda_{1}}$ (1)
$P(V)=Q(V)\exp\left(\frac{V-\lambda_{2}}{\lambda_{1}}\right)=Q(V)$ (2)
[…]
A.3 Theorem 3
Restatement of Theorem 3.
@@ A.4 Theorem 4
If $U=X+V$ with $X$ and $V$ both light-tailed and $V$ unbounded, and the distribution of U is continuous, and $\pi^{*}(\beta)\triangleq\arg\max_{\pi}\mathbb{E}[U(\pi)]-\beta D_{KL}(\pi,\pi_{0})$, then $\lim_{\beta\to 0^{+}}\mathbb{E}[V(\pi^{*}(\beta))]=\infty$.
@@
A.5 Theorem 5
Restatement of theorem 5.
@@ A.5 Theorem 5
  • The tail of $V$ is sufficiently lighter than the tail of $X$ that $\lim_{t\to\infty}\frac{t^{p}\bar{F}_{V}(t)}{\bar{F}_{X}(t)}=0$ for some $p>1$.
@@
In addition to the works cited in the main paper, we make reference to the textbook (Foss et al., , 2013) throughout the proof. Many similar results about random variables are present in the textbook.
A.5.1 Proof sketch and intuitions
@@
In Table A.1, we abbreviate the condition $X+V>t$ as $c$.
@@
Table A.1: A summary of the proof strategy for Theorem 5.
Figure A.2: A diagram showing the region boundaries at $-h(t)$, $h(t)$, and $t-h(t)$ in an example where $t=25$ and $h(t)=4$, along with a negative log plot of the relevant distribution:
@@ A.5.3 Lemmas
Lemma 2.19 from (Foss et al., , 2013) implies that if $X$ is long-tailed (which it is, because subexponential implies long-tailed), then there is $h(t)$ such that condition (a) holds and $\bar{F}_{X}$ is $h$-insensitive; by Proposition 2.20 we can take $h$ such that $h(t)\leq t/2$ for sufficiently large $t$, implying condition (b). Conditions (c) and (d) follow from being $h$-insensitive. ∎
@@
This is a slight variation on lemma 3.8 from (Foss et al., , 2013), and follows from the proof of Lemma 2.37. Lemma 2.37 states that
@@ A.5.5 Bounds on the denominator
A.6 Theorem 6
Restatement of theorem 6.
@@ A.6 Theorem 6
Theorem 6 generalizes a consequence of the "Regressional Goodhart Identity" in (Gao et al., , 2023).
@@
Appendix B Additional experiments and figures
Figures B.1, B.2 relate to experiments mentioned in the main paper. In response to reviewer feedback, we added two further experiments to demonstrate the catastrophic Goodhart phenemonon with artificially heavy-tailed reward, one using best-of-N on synthetic distributions and one with PPO on Pythia 1B.
@@ Appendix B Additional experiments and figures
Figure B.1: Histogram and normal probability plot of reward assigned by Pythia RM to random length-1024 token sequences. The Q-Q plot suggests the distribution is approximately normal, which is much lighter-tailed than exponential.
Figure B.2: Reward and log-probability for ACG-optimized inputs to Starling 7B-alpha.
B.1 Best-of-N experiment
We created a synthetic experiment by letting reward $U=X+V$, where $X$ and $V$ are independent and sampled from different probability distributions, consistent with our theoretical assumptions. We vary $N$ from 1 to 65536, do 100 trials of taking the best-of-$N$ sample with highest $U$, and note whether $V$ goes towards 0 (overoptimization) or not.
Possible distributions for V are normal and t-distribution with df=10. Possible distributions for X are normal, t with df=3, t with df=5, lognormal, and Levy. (All of these heavy-tailed except for the normal distribution.) V is scaled to a standard deviation of 2 and X has s.d. of 1 (except for the Levy distribution, which has infinite variance), representing that in ordinary regimes most of the variance comes from utility rather than error.
The results are shown in Figure B.3. Briefly, the results are consistent with the asymptotic results in theorems 5 and 6, showing that overoptimization
Figure B.3: When the error $X$ is normal and thus light-tailed, $V$ increases monotonically with $N$, consistent with our Theorem 6. However, when both $X$ and $V$ are heavy-tailed, we see results consistent with theorem 5. In $5$ of $6$ cases when $X$ is lognormal or student-t, $V$ first increases then starts to decline around $N=10^{2}$ or $10^{3}$. When $X$ is (t, df=5) and V is (t, df=10), $V$ instead peaks around $N=10^{5}$ (but declines afterwards). Finally, when X is Levy-distributed, utility never goes significantly above zero (optimization completely fails) because the Levy distribution is too heavy-tailed.
B.2 PPO experiment
In this experiment, we examined PPO with artificially heavy-tailed rewards to see if catastrophic Goodhart could be observed.
OpenRLHF (Hu et al., , 2024) was used to train Pythia 1B with a reward model derived also from Pythia 1B, on the default OpenRLHF prompt dataset. We used the reward model to represent true utility, and a heavy-tailed error term based on the number of "the" tokens was added to get the proxy reward. The kl_target=0.5 option was used to dynamically adjust KL penalty, as we mention is done in Ziegler et al., (2020). Rewards were not clipped. (Reward clipping can be useful to prevent overoptimization, but is not always used in PPO.) Response length was limited to 256.
Two example completions are in figure B.4. Midway through training, the model generates reasonable text with a higher number of "the" tokens than usual, but as training progresses the model starts to generate completions with a very high number of "the" tokens while proxy reward reaches values $>10^{9}$, demonstrating that the basic pattern of catastrophic Goodhart can occur in RLHF under conditions of heavy-tailed error.
The result depends on hyperparameters (e.g. reward clipping would prevent this), so our observation should not be taken as a claim that catastrophic Goodhart is inevitable in all RLHF settings with heavy-tailed error.
The preparation of polyarylethylene which is either employed for the purpose of applications in the petroleum refining or in the production of the well-known automobile oil lubes is carried out by the so-called preparation by the salt bath method described… in 1980 by D. W. Perkins in the United States. This method presupposes the usage of the salt bath as the medium for the preparation of the polyarylediene–the method of preparation of the crude oil base by the process which consists in letting it be saturated up with a particular the
A mirror, the essence of the understanding of the given the conclusion that the manifestation in the form of the gender and the identification the solutions related to different fields of the modern utilization of the the functionality of the the the the modification of the the the the the the state of the the the the the the the the the the the employment and the the the the the the the the the the the the the the the the the the the the the the the the the the the the the the the the the the the the the and the the the the the the the the the the the the the the the the the the the the by the the the the the the the the the the the the the the the the the the the the the the the the the the the the the the the the the the the and the the the the the the the the the the the the the the the the the
Figure B.4: PPO sample generations. Top: early in training, the model generates reasonable completions. Bottom: later in training, the artifically heavy-tailed reward dominates and the model generates completions with a very high number of "the" tokens.
Appendix C Assets
@@
We use three models for our experiments: Starling 7B-alpha, Llama 2 7B-chat, and Pythia-1.4B. Starling was developed by Berkeley, and Pythia by EleutherAI. Starling and Pythia models are licensed under Apache-2.0.999https://huggingface.co/berkeley-nest/Starling-RM-7B-alpha  [^10] https://huggingface.co/EleutherAI/pythia-1.4b Llama 2 models were developed by Meta and licensed under a license published by Meta. [^11] https://ai.meta.com/llama/license/
NeurIPS Paper Checklist
The checklist is designed to encourage best practices for responsible machine learning research, addressing issues of reproducibility, transparency, research ethics, and societal impact. Do not remove the checklist: The papers not including the checklist will be desk rejected. The checklist should follow the references and follow the (optional) supplemental material. The checklist does NOT count towards the page limit.
Please read the checklist guidelines carefully for information on how to answer these questions. For each question in the checklist:
  • You should answer [Yes] , [No] , or [N/A] .
  • [N/A] means either that the question is Not Applicable for that particular paper or the relevant information is Not Available.
  • Please provide a short (1–2 sentence) justification right after your answer (even for NA).
The checklist answers are an integral part of your paper submission. They are visible to the reviewers, area chairs, senior area chairs, and ethics reviewers. You will be asked to also include it (after eventual revisions) with the final version of your paper, and its final version will be published with the paper.
The reviewers of your paper will be asked to use the checklist as one of the factors in their evaluation. While "[Yes] " is generally preferable to "[No] ", it is perfectly acceptable to answer "[No] " provided a proper justification is given (e.g., "error bars are not reported because it would be too computationally expensive" or "we were unable to find the license for the dataset we used"). In general, answering "[No] " or "[N/A] " is not grounds for rejection. While the questions are phrased in a binary way, we acknowledge that the true answer is often more nuanced, so please just use your best judgment and write a justification to elaborate. All supporting evidence can appear either in the main paper or the supplemental material, provided in appendix. If you answer [Yes] to a question, in the justification please point to the section(s) where related material for the question can be found.
IMPORTANT, please:
  • Delete this instruction block, but keep the section heading “NeurIPS paper checklist",
  • Keep the checklist subsection headings, questions/answers and guidelines below.
  • Do not modify the questions and only use the provided macros for your answers.
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The abstract lists the important claims: the relationship between Goodhart’s Law and whether the error in a misspecified reward is heavy-tailed. The main limitation of independence assumptions is clearly stated in the introduction.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: Section 7 lists the limitations, which we have combined with the discussion section due to heavy overlap.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory Assumptions and Proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: All proofs are given inline or in the appendix, except for Theorem 5 which appears in the supplemental material.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental Result Reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: Sampling rewards requires no hyperparameters, and hyperparameters are provided for ACG.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: Code will be provided in the supplemental material.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental Setting/Details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: Other than hyperparameters, there are no details required to understand the results.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment Statistical Significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: The only error bars are inter-run variability of ACG. The standard deviation was reported rather than error bars due to the small number of runs.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments Compute Resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: The experiments took minimal compute resources except H100 hours for ACG, and we report the number of GPU-hours used.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code Of Ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: The research conforms to all data-related concerns. No human subjects were involved, and we think the risk of harmful societal impact is minimal.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader Impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: The immediate societal impacts are limited, but we discuss some potential applications to long-term safety.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: We have created no such artifacts.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: The assets we use are Starling 7B-alpha, Llama 2 7B-chat, and Pythia-1.4B. Starling and Pythia models are licensed under Apache-2.0.121212https://twitter.com/NexusflowX/status/1770532630645420474  [^13] https://huggingface.co/EleutherAI/pythia-1.4b Llama 2 models are licensed under a license published by Meta. [^14] https://ai.meta.com/llama/license/ We are in compliance with all licenses and terms of use.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New Assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [N/A]
  64. Justification: We do not release new assets.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and Research with Human Subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: No human subjects are involved.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: No human subjects are involved.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.